# The Cosine Schedule is Fisher-Rao-Optimal for Masked Discrete Diffusion Models

### v1

@@ 1 Introduction
In this work, we study the choice of discretisation schedule for masked discrete diffusion models from an information-geometric perspective (Amari, 2016). In particular, we show that the probabilistic structure of masked discrete diffusion allows for the closed-form computation of the Fisher-Rao metric of the induced probability path. In turn, this allows us to derive the optimal schedule - in terms of the Fisher-Rao geometry of the path - as the geodesic under this geometry, following the approach of Syed et al. (2022); Williams et al. (2024); Syed et al. (2024). Interestingly, we find that the optimal schedule recovers to the popularly-used cosine schedule (Nichol and Dhariwal, 2021).
@@ 3 Main Result
We see that this schedule coincides with the cosine schedule heuristically introduced in Nichol and Dhariwal (2021) (with $s=0$) for continuous diffusion models. A similar result was proved by Santos and Lin (2023) which showed for continuous diffusion model, when the data distribution $q_{0}(x_{0})$ is a Dirac measure - i.e. $q_{0}(x_{0})=\delta_{x}(x_{0})$, the cosine schedule is also optimal under the Fisher-Rao geometry. These results provide an interesting theoretical justification for the use of the cosine schedule.
@@ 4 Conclusion
We have shown that the optimal schedule under the Fisher-Rao geometry for masked discrete diffusion models recovers the popularly-used cosine schedule; we leave to future work the empirical validation of this schedule choice. Moreover, we hope this spurs further investigation into the information-geometric properties of generative models; for instance, considering optimal schedules under different metric choices. Finally, one limitation of our work is that we only consider the properties of the true probability path $q_{t}$. This does not take into account either the errors in the approximation of $q_{t}$ or the sampling scheme (Kim et al., 2025) used during generation. It is an interesting future direction to study the information geometry of these settings.
@@
LZ thanks Saifuddin Syed for introducing the information-geometric perspective on schedule tuning and Peter Potaptchik for helpful conversations. LZ is supported by the EPSRC CDT in Modern Statistics and Statistical Machine Learning (EP/S023151/1).

### v2

@@ 1 Introduction
In this work, we study the choice of discretisation schedule for masked discrete diffusion models from an information-geometric perspective (Amari, 2016). In particular, we show that the probabilistic structure of masked discrete diffusion allows for the closed-form computation of the Fisher-Rao metric of the induced probability path. In turn, this allows us to derive the optimal schedule—in terms of the Fisher-Rao geometry of the path—as the geodesic under this geometry, following the approach of Syed et al. (2022); Williams et al. (2024); Syed et al. (2024). Interestingly, we find that the optimal schedule recovers to the popularly-used cosine schedule (Nichol and Dhariwal, 2021).
@@ 3 Main Result
We see that this schedule coincides with the cosine schedule heuristically introduced in Nichol and Dhariwal (2021) (with $s=0$) for continuous diffusion models. A similar result was proved by Santos and Lin (2023) which showed for continuous diffusion model, when the data distribution $q_{0}(x_{0})$ is a Dirac measure—i.e. $q_{0}(x_{0})=\delta_{x}(x_{0})$, the cosine schedule is also optimal under the Fisher-Rao geometry. Another related result can be found in Theorem 6 from Chehab et al. (2023) which shows that the Fisher-Rao geodesic between some distributions $p_{0}$ and $p_{1}$ converges to the arithmetic annealing path with the cosine schedule, as the support of $p_{0},p_{1}$ become more disjoint. These results provide an interesting theoretical justification for the use of the cosine schedule.
@@ 4 Conclusion
We have shown that the optimal schedule under the Fisher-Rao geometry for masked discrete diffusion models recovers the popularly-used cosine schedule; we leave to future work the empirical validation of this schedule choice. Moreover, we hope this spurs further investigation into the information-geometric properties of generative models, for instance, considering optimal schedules under different metric choices. Finally, one limitation of our work is that we only consider the properties of the true probability path $q_{t}$. This does not take into account either the errors in the approximation of $q_{t}$ or the sampling scheme (Kim et al., 2025) used during generation. It is an interesting future direction to study the information geometry of these settings.
@@
LZ and SS thank Peter Potaptchik for helpful conversations. LZ is supported by the EPSRC CDT in Modern Statistics and Statistical Machine Learning (EP/S023151/1). SS acknowledges support from the NSERC Postdoctoral Fellowship.